# Sorting the Babble in Babel: Assessing the Performance of Language Detection Algorithms on the OpenAlex Database

### v1

@@ Abstract
Following a recent study on the quality of OpenAlex linguistic metadata (Céspedes et al.,, 2025), the present paper aims to optimize the latter through the design, use, and evaluation of various linguistic classification procedures based on the latest and most efficient automatic language detection algorithms. Starting from a multilingual set of manually-annotated samples of articles indexed in the database, different classification procedures are then designed, based on the application of a set of language detection algorithms on a series of corpora generated from different combinations of textual metadata of indexed articles. At sample level first, the performance of these different procedures for each of the main languages in the database is evaluated in terms of precision, recall, and processing time. Then, overall procedure performance is estimated at the database level by means of a probabilistic simulation of harmonically aggregated and weighted scores. Results show that procedure performance strongly depends on the importance given to each of the measures implemented: for contexts where precision is preferred, using the LangID algorithm on article titles, abstracts as well as journal names gives the best results; however, for all cases where recall is considered at least slightly more important than precision or as soon as processing times are given any kind of consideration, use of the FastSpell algorithm on article titles only outperforms all other alternatives. Given the lack of truly multilingual, large-scale bibliographic databases, it is hoped that these results help confirm and foster the unparalleled potential of the OpenAlex database for cross-linguistic, bibliometric-based research and analysis.
Keywords— OpenAlex, Automatic Language Detection, Bibliographic Databases, Multilingual Databases, Information Retrieval, Algorithm Evaluation, Natural Language Processing
@@ 1 Introduction
At the crossroads of linguistics and computer science research, monolingual automatic language detection is nowadays considered a solved task in the literature, with state-of-the-art models achieving above human-level performance in as much as 1,366 languages (Caswell et al.,, 2020). Integrated in various real-world applications, from translation services to user support and social media monitoring, language detection algorithms can also greatly enhance the organization, accessibility, and leveraging of information stored in databases. More precisely, by automatically identifying and categorizing the language of text attributes, these algorithms facilitate the sorting and indexing of multilingual data, which makes it easier to manage large, world-wide collections. Used in conjunction with machine translation algorithms, they can also support effective cross-language searches by allowing users to find pertinent information regardless of the language in which either the query or the relevant documents were written. Last but not least, language detection can benefit both data and resource management by respectively improving cross-linguistic data consistency and providing insights based on language distributions within the database.
@@
Despite these advantages, the use of language detection algorithms in bibliographic databases remains as rare as scarcely documented. The main reason for this gap in implementation and research is both historical and infrastructural, as bibliometric databases such as Scopus or the Web of Science have traditionally and quasi-exclusively focused on publications written in English (Ammon,, 2006; Archambault et al.,, 2005; Mongeon and Paul-Hus,, 2016; Demeter,, 2020; Vera-Baceta et al.,, 2019). This linguistic bias, while greatly reducing the need for multi- and cross-linguistic functionalities, also impacts the quality and scope of both research evaluation and research on research: being mostly blind to anything that eludes their monolingual lens, traditional bibliometric databases thus provide a fundamentally skewed view of the disciplinary and geographical landscape of the global scientific community (Beigel and Digiampietri,, 2022), which as a consequence distorts every initiative or investigation it informs and whose repercussions affect the neglected communities in various ways and degrees (Tennant,, 2020; Salatino and Ruiz,, 2021; Finardi,, 2022; Navarro et al.,, 2022).
@@
Recent open-source alternative to the main traditional bibliometric databases, OpenAlex (OpAl) stands out from its predecessors by offering a broader landscape of scholarly communication (Priem et al.,, 2022). Built on an open codebase and from a variety of bibliometric projects, services, databases, and repositories, it has been shown to outperform both the Web of Science and Scopus in terms of document and journal frequency (Alperin et al.,, 2024; Culbert et al.,, 2024; van Bellen et al.,, 2024; Jiao et al.,, 2023). Of particular relevance here, recent research has shown that OpAl also has a higher proportion of articles written in non-English languages than the two above-mentioned bibliographic databases (Céspedes et al.,, 2025; Vera-Baceta et al.,, 2019). In light of these considerations, not only does OpAl have the potential to offer a broader, multilingual, and as such more representative view of scholarly production, but its stronger allophone content makes it also more susceptible to benefit from the functionalities and advantages language detection algorithms can offer to multilingual databases.
@@
In that respect, while a quality assessment of OpAl language metadata has already been carried out (Céspedes et al.,, 2025), no attempt has yet been made yet to optimize it, that is, to find the best automatic language classification method for the articles it contains and indexes, based on its current metadata content and configuration. Such comparative assessment of language detection procedures will be the focus of the present paper. More precisely, the objective here is to apply different algorithms on various article metadata corpora extracted from a manually-annotated sample of OpAl-indexed publications in order to find the algorithm-corpus combination that performs best over a selected set of evaluation measures. The following section gives a detailed account of the methodological outline designed for this project, followed respectively by an analysis and discussion of the results obtained.
@@ 2.1 Sample Collection
The ground truth for this study is based on the annotation work carried out by Céspedes et al., (2025) in their assessment of language metadata quality in OpAl. For that paper, a multilingual sample of indexed articles was manually annotated following a two step process. First, two samples of indexed articles were randomly drawn from the database, a first one containing 50 articles for each of the 55 languages present in the database (the resulting sample included only $2701$ articles, as some languages had less than 50 indexed articles) and a second containing 300 documents for each of the 11 most frequent languages in the database. Then, each article sampled was manually annotated by not only inspecting its title and abstract, but also by accessing the full text of the corresponding document, identify its language, and then determine if that paper indeed corresponds to a scientific paper; entries whose documents couldn’t be accessed were ignored.

### v2

@@ Abstract
This project aims to compare various language classification procedures, procedures combining various Python language detection algorithms and metadata-based corpora extracted from manually-annotated articles sampled from the OpenAlex database. Following an analysis of precision and recall performance for each algorithm, corpus, and language as well as of processing speeds recorded for each algorithm and corpus type, overall procedure performance at the database level was simulated using probabilistic confusion matrices for each algorithm, corpus, and language as well as a probabilistic model of relative article language frequencies for the whole OpenAlex database. Results show that procedure performance strongly depends on the importance given to each of the measures implemented: for contexts where precision is preferred, using the LangID algorithm on the greedy corpus gives the best results; however, for all cases where recall is considered at least slightly more important than precision or as soon as processing times are given any kind of consideration, the procedure combining the FastSpell algorithm and the Titles corpus outperforms all other alternatives. Given the lack of truly multilingual, large-scale bibliographic databases, it is hoped that these results help confirm and foster the unparalleled potential of the OpenAlex database for cross-linguistic, bibliometric-based research and analysis.
Keywords— Bibliographic Databases, Automatic Classification, Natural Language Processing
@@ 1 Introduction
At the crossroads of linguistics and computer science research, monolingual automatic language detection is nowadays considered a solved task in the literature, with state-of-the-art models achieving above human-level performance in as much as 1,366 languages (Caswell et al.,, 2020). Integrated in various real-world applications, language detection algorithms can also greatly enhance the organization, accessibility, and leveraging of information stored in databases. More precisely, by automatically identifying and categorizing the language of text attributes, these algorithms facilitate the sorting and indexing of multilingual data, which makes it easier to manage large, world-wide collections. Used in conjunction with machine translation algorithms, they can also support effective cross-language searches by allowing users to find pertinent information regardless of the language in which either the query or the relevant documents were written. Last but not least, language detection can benefit both data and resource management by respectively improving cross-linguistic data consistency and providing insights based on language distributions within the database.
@@
Despite these advantages, the use of language detection algorithms in bibliographic databases remains as rare as scarcely documented. The main reason for this gap in implementation and research is both historical and infrastructural, as databases such as Scopus or the Web of Science have traditionally and quasi-exclusively focused on publications written in English (Ammon,, 2006; Archambault et al.,, 2005; Mongeon and Paul-Hus,, 2016; Demeter,, 2020; Vera-Baceta et al.,, 2019). This linguistic bias, while greatly reducing the need for multi- and cross-linguistic functionalities, also impacts the quality and scope of both research evaluation and research on research: being mostly blind to anything that eludes their monolingual lens, traditional bibliometric databases thus provide a fundamentally skewed view of the disciplinary and geographical landscape of the global scientific community (Beigel and Digiampietri,, 2022), which as a consequence distorts every initiative or investigation it informs and whose repercussions affect the neglected communities in various ways and degrees (Tennant,, 2020; Salatino and Ruiz,, 2021; Finardi,, 2022; Navarro et al.,, 2022).
@@
Recent open-source alternative to the main traditional bibliometric databases, OpenAlex (OpAl) stands out from its predecessors by offering a broader landscape of scholarly communication (Priem et al.,, 2022). Built on an open codebase and from a variety of bibliometric projects, services, databases, and repositories, it has been shown to outperform both the Web of Science and Scopus in terms of document and journal frequency (Alperin et al.,, 2024; Culbert et al.,, 2024; van Bellen et al.,, 2024; Jiao et al.,, 2023). Of particular relevance here, recent research has shown that OpAl also has a higher proportion of articles written in non-English languages than the two above-mentioned bibliographic databases (Céspedes et al.,, 2025; Vera-Baceta et al.,, 2019). In light of these considerations, not only does OpAl have the potential to offer a broader, multilingual, and as such more representative view of scholarly production at the global scale, but its stronger allophone content makes it also more susceptible to benefit from the functionalities and advantages language detection algorithms can offer to multilingual databases.
Regarding that last point, previous studies have however shown that language detection algorithm performance in multilingual context can vary significantly depending on the languages involved or on the specific characteristics of the dataset. In that perspective, effective multilingual database management requires linguistic classification procedures that are not only tailored to the specific linguistic features and metadata characteristics of the dataset, but also based on algorithms that are robust to noise in the data attributes considered (Balazevic et al.,, 2016; Polasi and Sri Rama Krishna,, 2016; Sujaini and Negara,, 2024). In the specific case of OpenAlex, while a quality assessment of OpAl language metadata has already been carried out (Céspedes et al.,, 2025), no attempt has yet been made yet to optimize the quality of its language metadata, that is, to find the best linguistic classification procedure for the articles it contains and indexes, based on its current metadata content and configuration.
@@
Such comparative assessment of language detection procedures will be the focus of the present paper. More precisely, the objective here is to apply different algorithms on various article metadata corpora extracted from a manually-annotated sample of OpAl-indexed publications in order to find the algorithm-corpus combination that performs best over a selected set of evaluation measures. The following section gives a detailed account of the methodological outline designed for this project, followed respectively by an analysis and discussion of the results obtained.
@@ 2.1 Sample Collection
The ground truth for this study is based on the annotation work carried out by Céspedes et al., (2025) in their assessment of language metadata quality in OpAl. For that paper, a multilingual sample of indexed articles was manually annotated following a two step process. First, two samples of indexed articles were randomly drawn from the database, a first one containing 50 articles for each of the 55 languages present in the database (the resulting sample included only $2701$ articles, as some languages had less than 50 indexed articles) and a second containing 300 documents for each of the 11 most frequent languages in the database. Then, each article sampled was manually annotated by not only inspecting its title and abstract, but also by accessing the full text of the corresponding document, identify its language, and then determine if that paper indeed corresponds to a scientific paper; entries whose documents couldn’t be accessed were ignored (see (Céspedes et al.,, 2025) for more information about these samples).